# Modified Langevin noise formalism for multiple quantum emitters in dispersive electromagnetic environments

### v1

@@ VI.2 Entanglement Decay and Revivals
Since the early days of quantum information theory [28], quantum entanglement [29, 30] has been recognized as a key resource for a variety of tasks [31]. During the past two decades, its generation [25, 32, 33, 34, 35], sudden death [36, 37, 38, 39], birth [40], and degradation [41, 42, 43, 44] under multiple external environments [45] have been extensively investigated. In particular, if the quantum bipartite system undergoes non-Markovian dynamics [46, 47, 48, 49, 33, 35], it has been shown that the quantum state experiences revivals in its entanglement features. More recently, the generation of entangled states of pairs of quantum emitters using plasmonic [50], as well as photonic 1D waveguides [5] through bound states in the continuum [51, 52], has been studied.
@@
Allowing the medium- and scattering-assisted reservoirs to be in thermal states at different temperatures, with a the material-assisted reservoir playing the role of the hot bath, see Fig. 4, leads to increased populations of the excited states of the emitters. This behavior is justified by the increased probability of photon absorption from the warmer reservoir, which partially counteracts emission.
@@
The properties of the reduced states shown in Figs. 3 and 4 stem from the evolution of the initially pure entangled state of the composite system under its coupling to the photon reservoirs. As the system relaxes, the state becomes progressively mixed, while initial quantum correlations, such as entanglement, are progressively degraded. This process is illustrated in Fig. 5, where the entanglement negativity $\mathcal{N}(\hat{\rho}_{12})$ is plotted as a function of time for different temperatures of the thermal states of the electromagnetic environment and frequency detuning of the emitter transition frequencies. The detrimental impact of finite-temperature bath states on negativity is due to the combined effect of increased populations of excited emitter states and of the enhanced decay of two-body correlations, sustained by off-diagonal terms of $\hat{\rho}_{12}$. The Rabi-like interaction with the electromagnetic environment thus leads the composite state to a separable state $\mathcal{N}(\hat{\rho}_{12})=0$. This relaxation is governed by the interplay between direct- and cross- correlations, both of which extend over a broad frequency range due to the presence of the scattering-assisted reservoir.
@@
At stronger coupling, the physics becomes richer. Figure 6 shows the behavior of $\mathcal{N}(\rho_{12})$ for different values of the dipole momentum $\mu$, inverse temperatures of the electromagnetic environment, and fixed detuning of the emitter frequencies. Negativity undergoes sudden death at short times, followed by revivals of decreasing amplitudes at later times, hallmark of non-Markovian effects [35], arising from emitter-photon correlations, which are present even in the RWA limit [46]. However, in the case of Eq. (72), Hamiltonian terms beyond RWA can be proved to reduce the amplitudes of revivals. In addition, the finite-temperature states of the medium-assisted reservoir noticeably curb the entanglement revivals.
@@
The analysis of negativity is repeated in Fig. 7 for the configuration of Fig. 2, where two emitters interacts with two rods and a nano-disk, under equal, low temperatures of the medium and scattering-assisted reservoirs. Transient, short-time entanglement revivals exhibit a marked dependence on detuning: for zero detuning and moderate coupling, the revivals are enhanced at short times. Altogether, these results highlight how the detailed structure of the medium- and scattering-assisted spectral density matrices affects the degradation and revival of entanglement, suggesting concrete routes to optimize quantum correlations in structured nanophotonic environments.
@@
Figure 3: Spherical particle, see Fig. 1. Relaxation dynamics of populations of quantum emitters: $p_{a}^{e}(t)$ is the population of the quantum emitter $a$, for $a=1,2$, in the excited state $|e\rangle$. The initial composite state has been chosen as $|\hat{\rho}_{12}(0)=\outerproduct{\Psi^{-}}{\Psi^{-}}$, while $\mu=10^{-3}\omega_{p}$ and $\beta_{S}=\beta_{M}=1000$. solid blue (red) curves denote the case $\Omega_{1}=0.54\omega_{p},\Omega_{2}=\omega_{p}$, while cyan (orange) are computed for $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$.
@@
Figure 4: Spherical particle, see Fig. 1. Relaxation dynamics of populations of the quantum emitters: $p_{a}^{e}(t)$ is the population of the quantum emitter $a$, for $a=1,2$, in the excited state $|e\rangle$. The initial composite state has been chosen as $|\rho_{12}(0)=\outerproduct{\Psi^{-}}{\Psi^{-}}$, while $\mu=10^{-3}\omega_{p}$ and $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$. Solid blue (red) curves are computed for $\beta_{S}=\beta_{M}=1000$, and dashed curves denote the case $\beta_{M}=1,\beta_{S}=1000$.
@@
Figure 5: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ with $\mu=10^{-3}\omega_{p}$, corresponding to the same initial state as in Figs. 3 and 4. Solid blue(red) lines denote the cases $\Omega_{1}=0.54\omega_{p},\Omega_{2}=\omega_{p}$ and $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$ with equal inverse temperatures $\beta_{S}=\beta_{M}=1000$, respectively, while dashed curves denote the same frequency choices for $\beta_{M}=1$ and $\beta_{S}=1000$.
@@
Figure 6: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$, computed for $\Omega_{1}=0.54\omega_{p},\Omega_{2}=\omega_{p}$ and the same initial state as in Figs.3 and 4. Solid lines denote the cases $\mu=0.01,0.02,0.04$ and inverse temperatures $\beta_{S}=\beta_{M}=1000$, while dashed lines denote analogous curves computed for $\beta_{M}=1,\beta_{S}=1000$.
@@
Figure 7: Rod-nanodisk configuration, see Fig. 2. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$, computed for $\beta_{M}=\beta_{S}=1000$ and the same initial state as in Figs.3 and 4, for different choices of transition frequencies and dipole momenta.
@@ VI.3 Entanglement Generation
We now address the creation of entanglement starting from separable state preparations. In particular, we confine the initial state of the two emitters to convex mixtures of product states $\sum_{k}p_{k}\,\rho^{k}_{1}\otimes\rho^{k}_{2}$, and focus on entanglement generation in the course of relaxation, keeping in mind that separable states may still display quantum correlations beyond entanglement [58]. To isolate environment-induced effects, we consider the pure product $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, so that no initial classical or quantum correlations are present. The electromagnetic environment is associated with the dielectric sphere of Fig. 1; where we set the detuning of the two emitter frequencies to zero, i.e., $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$.
@@
In Fig. 8, we first study the evolution of the composite state for fixed coupling strengths and inverse temperatures. The results show that for sufficiently long times, $\mathcal{N}(\hat{\rho}_{12})$ turns positive and starts increasing as a function of time, i.e., a degree of entanglement between the two emitters can be generated as a result of evolution. This property can be traced back to the action of the cross-coupling terms in Eq. (72), which allow for photon-mediated interactions between the two emitters. Similarly to Sect. VI.2, in the case of finite temperature of the medium-assisted reservoir, the system evolves in a separable state for very long times, i.e., in the considered setting, finite temperatures turn out to be detrimental for mediated interactions.
@@
A further dynamical signature of the influence of the cross-coupling interactions along with their spectral features is reported in Fig. 9, where we compare the negativity $\mathcal{N}(\hat{\rho}_{12})$ for time-evolved states of identical emitters with $\Omega_{1}=\Omega_{2}$ while parametrically varying the common transition frequency. When the transition frequencies are tuned near the peak of the cross spectral density $\mathcal{J}_{12}(\omega)$, the value of negativity at moderately long times can be enhanced.
@@
It is also natural to ask whether or not the generation of entanglement states of the emitters at long times decreases with the coupling strength. In principle, increasing the coupling strength of the two emitters can lead to enhancements of these long-time quantum correlations. However, due to the form of Eq.(72), relaxation of the individual emitter states would also increase, as well as decoherence in the state of the composite system. As a consequence, competing effects take place in the course of the dynamics, resulting in a non-monotonic behavior of entanglement with the magnitude of $\mu$. The effects of this competition are evident in Fig. 10, where we plot $\mathcal{N}(\hat{\rho}_{12})$ at low environment temperatures for increasing $\mu$. Notice that the onset time of nonzero negativity increases with $\mu$. This hints at possible steady states that can sustain entanglement between the emitters. Establishing the existence and nature of such steady states would require different numerical approaches [59], which is left for future work.
@@
Figure 8: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ for the initial composite state $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$, $\mu=10^{-3}\omega_{p}$, computed for $\beta_{S}=\beta_{M}=1000$ (solid line) and $\beta_{M}=1,\beta_{S}=1000$ (dashed line).
@@
Figure 9: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ for the initial composite state $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, $\mu=10^{-3}\omega_{p}$, $\beta_{S}=\beta_{M}=1000$, computed for different values of the emitter frequencies $\Omega_{1}=\Omega_{2}$ in the range $0.30\omega_{p},0.80\omega_{p}$].
@@
Figure 10: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ for the initial composite state $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$$\beta_{S}=\beta_{M}=1000$, computed for different dipole strengths $\mu$ in the range $1\cdot 10^{-3}\omega_{p},5\cdot 10^{-3}\omega_{p}]$.
@@ VII Conclusions and outlook
Macroscopic quantum electrodynamics in the context of quantum nanophotonics is a robust framework for describing from an ab initio perspective how a collection of quantum emitters interacts with complex nanophotonic structures, which can be accounted for by using macroscopic Maxwell equations. In this article, we have introduced an extension of the modified Langevin noise approach we recently developed for a single quantum emitter [1], [2] to the case of several quantum emitters. In the modified Langevin noise formalism, the electromagnetic environment consists of two continuous bosonic reservoirs, a medium-assisted bosonic reservoir and a scattering-assisted bosonic reservoir that in general are initially in different quantum states. We have first demonstrated how to define and obtain the medium-assisted and scattering-assisted spectral densities, two distinct matrix-valued functions that fully characterize the interaction of the electromagnetic environment with the quantum emitters. These matrices cannot be evaluated from the knowledge of the dyadic Green function of the dielectric objects at the positions of the quantum emitters. We have shown that they can be effectively computed numerically by surface integral equation formulations of linear classical electromagnetic scattering. We have then introduced the temperature-dependent medium and temperature-dependent scattering-assisted spectral density matrices to describe electromagnetic environments with medium-assisted and scattering-assisted bosonic reservoirs initially in thermal quantum states at different temperatures. Eventually, we have demonstrated that the time evolution of the reduced density operator of the quantum emitters can be obtained by introducing a surrogate bosonic environment, initially in a vacuum quantum state, with an effective spectral density matrix given by the sum of the temperature-dependent medium and temperature-dependent scattering-assisted spectral density matrices. Once the effective spectral density matrix has been calculated, the reduced dynamics of the quantum emitters can be evaluated by using standard approaches for non-Markovian open quantum systems. It is important to note that this framework is applicable not only for deriving precise solutions in quantum nanophotonics, but also as a basis for developing simplified models.
We apply the approach to two quantum emitters that interact with a metallic spherical particle and a metallic nanostructure composed of two rods and a disk. In particular, we found entanglement decay and revivals, as well as entanglement generation. These results underline the utmost importance of the detailed structure of the environment spectral densities in the degradation of quantum correlations. This suggests that the proposed approach can be used to optimize genuinely quantum features of quantum emitters in structured nanophotonics environments.
Future work should focus on extending the proposed approach to the study of molecular polaritons in a dilute ensemble of molecular quantum emitters in optical resonators.

### v2

@@ VI.2 Entanglement Decay and Revivals
Since the early days of quantum information theory [35], quantum entanglement [36, 37] has been recognized as a key resource for a variety of tasks [38]. During the past two decades, its generation [32, 39, 40, 41, 42], sudden death [43, 44, 45, 46], birth [47], and degradation [48, 49, 50, 51] under multiple external environments [52] have been extensively investigated. In particular, if the quantum bipartite system undergoes non-Markovian dynamics [53, 54, 55, 56, 40, 42], it has been shown that the quantum state experiences revivals in its entanglement features. More recently, the generation of entangled states of pairs of quantum emitters using plasmonic [18, 57], as well as photonic 1D waveguides [5] through bound states in the continuum [58, 59], has been studied.
@@
Allowing medium- and scattering-assisted reservoirs to be initially in thermal states at different temperatures, with the medium-assisted reservoir playing the role of the hot bath (Fig. 4) yields larger excited-state populations for both emitters. This behavior is explained by the increased probability of photon absorption from the warmer reservoir, which partially counteracts emission.
@@
The reduced-state dynamics in Figs. 3 and 4 stem from the evolution of the initially pure entangled state of the composite system under coupling to the photon reservoirs. As the system relaxes, the state becomes progressively mixed, while initial quantum correlations, such as entanglement, are progressively degraded. This process is illustrated in Fig. 5, where the entanglement negativity $\mathcal{N}(\hat{\rho}_{12})$ is plotted as a function of time for different temperatures of the thermal states of the electromagnetic environment and for different detuning of the emitter transition frequencies. The detrimental impact of finite-temperature bath states on negativity is due to the combined effect of increased populations of excited emitter states and of the enhanced decay of two-body correlations, sustained by off-diagonal terms of $\hat{\rho}_{12}$. The Rabi-like interaction with the electromagnetic environment thus leads the composite state to a separable state $\mathcal{N}(\hat{\rho}_{12})=0$. This relaxation is governed by the interplay between direct- and cross-correlations, both of which extend over a broad frequency range due to the presence of the scattering-assisted reservoir.
@@
With stronger coupling, the dynamics becomes richer. Figure 6 shows the behavior of $\mathcal{N}(\hat{\rho}_{12})$ for different values of the dipole momentum $\mu$, inverse temperatures, and fixed detuning of the emitter frequencies. Negativity undergoes sudden death at short times, followed by revivals of decreasing amplitudes at later times, a hallmark of non-Markovian effects [42], arising from emitter-photon correlations, which are present even in the RWA limit [53]. In the present model, counter-rotating terms in Eq. (70) tend to suppress revival amplitudes compared to a purely RWA description. In addition, the finite-temperature states of the medium-assisted reservoir noticeably curb the entanglement revivals.
@@
The analysis of negativity is repeated in Fig. 7 for the configuration of Fig. 2, where two emitters interact with two rods and a nano-disk, at equal, low temperatures of the medium and scattering-assisted reservoirs. Transient, short time entanglement revivals exhibit a marked dependence on detuning: for zero detuning and moderate coupling, the revivals are enhanced at short times. Altogether, these results highlight how the detailed structure of the medium- and scattering-assisted spectral density matrices affects the degradation and revival of entanglement, suggesting concrete routes to optimize quantum correlations in structured nanophotonic environments.
@@
Figure 3: Spherical particle, see Fig. 1. Relaxation dynamics of the excited-state populations $p_{e}^{a}(t)$ for the two quantum emitters ($a=1,2$). The initial composite state is $|\hat{\rho}_{12}(0)=\outerproduct{\Psi^{-}}{\Psi^{-}}$, the parameters are $\mu=10^{-3}\omega_{p}$ and $\beta_{S}=\beta_{M}=1000$. Blue (red) curves refer to emitter $a=1$ ($a=2$) for $\Omega_{1}=0.54\omega_{p},\Omega_{2}=\omega_{p}$, while cyan (orange) curves refer to emitter $a=1$ ($a=2$) for $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$ for emitter $a=1$ ($a=2$). The cutoff frequency of the environment is set to $\omega_{\text{cut}}=4.0\,\omega_{p}$.
@@
Figure 4: Spherical particle, see Fig. 1. Relaxation dynamics of the excited-state populations $p_{e}^{a}(t)$ for the two quantum emitters ($a=1,2$). The initial composite state is set to $|\rho_{12}(0)=\outerproduct{\Psi^{-}}{\Psi^{-}}$, while $\mu=10^{-3}\omega_{p}$ and $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$. Solid curves are correspond to $\beta_{S}=\beta_{M}=1000$, whereas dashed curves illustrate the case $\beta_{M}=1,\beta_{S}=1000$. The cutoff frequency of the environment is set to $\omega_{\text{cut}}=4.0\omega_{p}$.
@@
Figure 5: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ with $\mu=10^{-3}\omega_{p}$, corresponding to the same initial state as in Figs. 3 and 4. Solid lines denote the cases $\Omega_{1}=0.54\omega_{p},\Omega_{2}=\omega_{p}$ and $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$ with equal inverse temperatures $\beta_{S}=\beta_{M}=1000$, respectively, while dashed curves denote the same frequency choices for $\beta_{M}=1$ and $\beta_{S}=1000$. The cutoff frequency of the environment is set to $\omega_{\text{cut}}=4.0\omega_{p}$.
@@
Figure 6: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$, computed for $\Omega_{1}=0.54\omega_{p},\Omega_{2}=\omega_{p}$ and the same initial state as in Figs. 3 and 4. Solid lines denote the cases $\mu=0.01,0.02,0.04$ and inverse temperatures $\beta_{S}=\beta_{M}=1000$, while dashed lines denote analogous curves computed for $\beta_{M}=1,\beta_{S}=1000$. The cutoff frequency of the environment is set to $\omega_{\text{cut}}=4.0\omega_{p}$.
@@
Figure 7: Rod-nanodisk configuration, see Fig. 2. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$, computed for $\beta_{M}=\beta_{S}=1000$ and the same initial state as in Figs.3 and 4, for different choices of transition frequencies and dipole momenta. The cutoff frequency of the environment is set to $\omega_{\text{cut}}=4.0\omega_{p}$.
@@ VI.3 Entanglement Generation
We now address the creation of entanglement starting from separable state preparations. In particular, we consider initial two-emitter states that are convex mixtures of product states $\sum_{k}p_{k}\,\hat{\rho}^{k}_{1}\otimes\hat{\rho}^{k}_{2}$, and examine whether entanglement emerges in the course of relaxation, keeping in mind that separable states may still display quantum correlations beyond entanglement [65]. To isolate environment-induced effects, we take the initial two-emitter state to be the pure product $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, so that no initial classical or quantum correlations are present. The electromagnetic environment is associated with the dielectric sphere of Fig. 1; where we set the detuning of the two emitter frequencies to zero, i.e., $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$.
@@
Figure 8 shows the evolution of $\mathcal{N}(\hat{\rho}_{12})$ for fixed coupling strengths and inverse temperatures. At sufficiently long times, the negativity $\mathcal{N}(\hat{\rho}_{12})$ becomes positive and then grows, demonstrating that entanglement between the two emitters generated dynamically from an initially separable state. This behavior originates from the cross-coupling terms in Eq. (70), which allow photon-mediated interactions between the two emitters. Similarly to Sec.VI.2, in the case of finite temperature of the medium-assisted reservoir, the system evolves in a separable state for very long times, i.e., in the considered setting, finite temperatures turn out to be detrimental for mediated interactions.
@@
An additional dynamical signature of the influence of the cross-coupling interactions along with their spectral features is reported in Fig. 9, where we compare the negativity $\mathcal{N}(\hat{\rho}_{12})$ for time-evolved states of identical emitters with $\Omega_{1}=\Omega_{2}$ for different values of the common transition frequency. When the transition frequencies are tuned close to the peak of the cross spectral density $\mathcal{J}_{12}(\omega)$, the value of negativity at moderately long times can be enhanced.
@@
It is also natural to ask how the generation of entanglement at long times depends on the coupling strength. In principle, increasing the coupling strength of the two emitters can enhance the build-up of long-time quantum correlations. However, due to the form of Eq. (70), relaxation of the individual emitter states would also increase, as well as decoherence in the state of the composite system. The resulting competition leads to a non-monotonic dependence of the generated entanglement on $\mu$. This behavior is evident in Fig. 10, where we plot $\mathcal{N}(\hat{\rho}_{12})$ at low environment temperatures for increasing $\mu$. As $\mu$ grows, the onset of nonzero negativity is delayed. For sufficiently strong coupling, it is evident that the negativity undergoes a transient time behavior: it reaches a local maximum at intermediate times and then decays, approaching zero at long times. Overall, the data in Fig. 10 point towards the existence of steady states that cannot sustain entanglement between the emitters, at least for moderate coupling strengths.
However, establishing the existence and nature of such steady states would require different numerical approaches [66], which is left for future work.
@@
Figure 8: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ for the initial composite state $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$, $\mu=10^{-3}\omega_{p}$, computed for $\beta_{S}=\beta_{M}=1000$ (solid line) and $\beta_{M}=1,\beta_{S}=1000$ (dashed line). The cutoff frequency of the environment is set to $\omega_{\text{cut}}=4.0\omega_{p}$.
@@
Figure 9: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ for the initial composite state $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, $\mu=10^{-3}\omega_{p}$, $\beta_{S}=\beta_{M}=1000$, computed for different values of the emitter frequencies $\Omega_{1}=\Omega_{2}$ in the range $0.30\omega_{p},0.80\omega_{p}]$. The cutoff frequency of the environment is set to $\omega_{\text{cut}}=4.0\omega_{p}$.
@@
Figure 10: Spherical particle, see Fig. 1. Time evolution of negativity $\mathcal{N}(\hat{\rho}_{12})$ for the initial composite state $|\hat{\rho}_{12}(0)=\outerproduct{eg}{eg}$, $\Omega_{1}=\Omega_{2}=0.54\omega_{p}$$\beta_{S}=\beta_{M}=1000$, computed for different dipole strengths $\mu$ in the range $5\cdot 10^{-4}\omega_{p},5\cdot 10^{-3}\omega_{p}]$. The cutoff frequency of the environment is set to $\omega_{\text{cut}}=4.0\omega_{p}$.
@@ VII Conclusions and outlook
Macroscopic quantum electrodynamics (QED) provides an efficient ab-initio framework for describing how quantum emitters interact with structured nanophotonic environments. In this work, we extended the modified Langevin noise approach we recently developed for a single quantum emitter [1, 2] to the case of multiple quantum emitters. Within this formalism, the electromagnetic environment is represented by two continuous bosonic reservoirs, a medium-assisted bosonic reservoir and a scattering-assisted bosonic reservoir, which may, in general, be prepared in different initial quantum states.
@@
We defined two matrix-valued spectral densities, the medium-assisted and scattering-assisted spectral density matrices, which fully characterize the interaction of the electromagnetic environment with the quantum emitters. Importantly, these matrices cannot be reconstructed from the sole knowledge of the dyadic Green function evaluated at the emitter positions. We showed that they can be computed efficiently within classical electromagnetics by means of surface-integral-equation formulations of linear scattering. We then introduced temperature-dependent medium and temperature-dependent scattering-assisted spectral density matrices to describe electromagnetic environments with medium-assisted and scattering-assisted bosonic reservoirs initially in thermal quantum states at different temperatures. Finally, we demonstrated that the time evolution of the reduced density operator of the quantum emitters can be obtained by introducing a surrogate bosonic environment, initially in a vacuum quantum state, whose effective spectral density matrix is the sum of the temperature-dependent medium and temperature-dependent scattering-assisted spectral density matrices. Once this effective spectral density matrix is known, the reduced dynamics of the quantum emitters can be evaluated with standard methods for non-Markovian open quantum systems.
We apply this approach to two quantum emitters coupled to a metallic spherical particle and a metallic nanostructure composed of two rods and a disk. In these representative geometries, we observe entanglement decay and pronounced revivals, as well as the emergence of entanglement from an initial product state. These findings highlight that the fate of quantum correlations is governed not merely by local decay rates, but by the full, frequency-dependent structure of the environment’s spectral-density matrices, which controls both the degradation and the generation of quantum correlations. More broadly, they indicate that the proposed approach provides a practical route to engineering, and ultimately optimizing, genuinely quantum features of emitter dynamics in structured nanophotonic environments driven out of thermal equilibrium.
A natural next step is to extend the framework to molecular polaritons, focusing on dilute ensembles of molecular emitters coupled to optical resonators.